Circular No. 8637

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 YJ_{35} (LINEAR) An apparently asteroidal object discovered by the LINEAR project

(discovery observation below from MPEC 2005-A47 and MPS 124152) has been found by S. S. Sheppard to have a centrally condensed coma with a tail ∼ 1′′.5 wide and ∼ 10′′ long in p.a. 330^{o} on CCD frames taken on Nov. 30.34 and Dec. 1.33 UT with the Dupont 2.5-m reflector at Las Campanas (MPEC 2005-X12; T = 2005 Mar. 3, q = 1.78 AU, i = 52^{o}, e = 0.9999).

2004 UT  α_{2000}  δ_{2000}  Mag.
Dec. 31.07192  21^{h}09^{m}18.^{s}11  +69^{o}18′19′′.5  19.4

COMET 101P/CHERNYKH E. J. Christensen has found an object with a well-condensed 15′′ coma

moving ∼ 21′.5 ahead of what has been presumed to be the primary component of comet 101P (corresponding to ΔT = −0.9 day) on 90-s images taken with the Catalina Sky Survey 0.68-m Schmidt telescope on Nov. 30.2 UT; his Dec. 2.2 images show a 15′′ asymmetric coma elongated in p.a. 50^{o} (the ∼ 2.5-mag-brighter component showing a 10′′ central condensation surrounded by a 1′ asymmetric coma also oriented in p.a. 50^{o}). Observations with the Mt. Lemmon Survey 1.5-m reflector by A. D. Grauer and E. C. Beshore on Dec. 1.2 show this fainter component to have a broad 10′′–15′′ tail in p.a. ∼ 57^{o}. The available astrometry is provided below. Following the reasoning by Sekanina (IAUC 5391), it is unlikely that the companion observed in 1991 (IAUC 5347) has survived.

2005 UT α_{2000} δ_{2000} Mag. Observer
Nov. 30.19707 0^{h}19^{m}57.^{s}46 −4^{o}57′00′′.8 17.8 Christensen
30.20586 0 19 57.82 −4 56 58.5 17.9  _{”}
30.21450 0 19 58.15 −4 56 55.7 18.3  _{”}
30.22309 0 19 58.44 −4 56 51.7 18.2  _{”}
Dec. 1.23064 0 20 36.93 −4 50 52.2 17.9 Grauer
1.23152 0 20 36.97 −4 50 51.8 18.1  _{”}
1.23241 0 20 36.99 −4 50 51.4 18.2  _{”}
1.23329 0 20 36.95 −4 50 51.9  _{”}
2.20865 0 21 15.62 −4 44 53.9 17.9 Christensen
2.21893 0 21 15.98 −4 44 49.8 18.3  _{”}
2.22922 0 21 16.45 −4 44 46.3 18.1  _{”}
2.23962 0 21 16.81 −4 44 43.0 17.8  _{”}
2005 December 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green